# Towards Open Vocabulary Learning: A Survey

### v1

@@ Abstract
In the field of visual scene understanding, deep neural networks have made impressive advancements in various core tasks like segmentation, tracking, and detection. However, most approaches operate on the close-set assumption, meaning that the model can only identify pre-defined categories that are present in the training set. Recently, open vocabulary settings were proposed due to the rapid progress of vision language pre-training. These new approaches seek to locate and recognize categories beyond the annotated label space. The open vocabulary approach is more general, practical, and effective compared to weakly supervised and zero-shot settings. This paper provides a thorough review of open vocabulary learning, summarizing and analyzing recent developments in the field. In particular, we begin by comparing it to related concepts such as zero-shot learning, open-set recognition, and out-of-distribution detection. Then, we review several closely related tasks in the case of segmentation and detection, including long-tail problems, few-shot, and zero-shot settings. For the method survey, we first present the basic knowledge of detection and segmentation in close-set as the preliminary knowledge. Next, we examine various scenarios in which open vocabulary learning is used, identifying common design elements and core ideas. Then, we compare the recent detection and segmentation approaches in commonly used datasets and benchmarks. Finally, we conclude with insights, issues, and discussions regarding future research directions. To our knowledge, this is the first comprehensive literature review of open vocabulary learning. We keep tracing related works at https://github.com/jianzongwu/Awesome-Open-Vocabulary.
@@ 1 Introduction
Deep neural networks have revolutionized scene understanding tasks, including object detection, segmentation and tracking [1, 2, 3, 4]. Nonetheless, using conventional approaches for real-world applications can be challenging due to limitations such as inadequate class annotations, close-set class definitions, and costly labeling expenses. These limitations can increase the difficulty and cost of implementing a deep model on new scenes, particularly when the number of categories or concepts in the scene is greater than what was included in the training dataset.
@@
For example, object detection is a core task in computer vision that involves scene understanding. It requires human annotations for each category and each object location, which can be costly and time-consuming. For instance, the COCO dataset [5], which is widely used for benchmarking object detection algorithms, only includes 80 categories. But in reality, natural scene images often have more than 80 different types of objects. To extend object detectors to cover all these categories, we would need to incur significant annotation costs. A key focus of current research is developing methods to train more flexible object detectors on the subset of COCO with base classes and let them identify new or unfamiliar objects without requiring additional annotations.
@@
Several previous solutions are to adopt zero-shot learning (ZSL) [6, 7, 8]. These approaches extend a detector to generalize from annotated (seen) object classes to other (unseen) categories. The annotations of seen object classes are used during the training, while the annotations of unseen classes are strictly unavailable during training. Most approaches adopt word embedding projection to constitute the classifier for unseen class classification.
@@
However, these approaches come with a number of limitations. ZSL, in particular, is highly restrictive. Typically, these methods lack examples of unseen objects and treat these objects as background object during training. As a result, during inference, the model identifies novel classes solely based on their pre-defined word embeddings [9, 10, 11], thereby limiting exploration of the visual information and relationships of those unseen classes. This is why ZSL approaches have been shown to yield unsatisfying results in novel classes.
@@
To handle the above issue and make the detector extendable, open vocabulary object detection [12, 13] is proposed. Also, open vocabulary segmentation is proposed [14, 15]. Similarly, the model trains on base classes and inferences on both base and novel classes. The key difference between zero-shot learning and open vocabulary is that one can use visual-related language vocabulary data like image captions as auxiliary supervision in open vocabulary settings. The motivations to use language data as auxiliary weak supervision are: 1) Language data requires less labeling effort and thus is more cost-friendly. The visual-related language data, like image captions, is widely available and more affordable to obtain compared to the box or mask annotations. 2) Language data provides a larger vocabulary size and thus is more extendable and general. For example, words in captions are not limited to the pre-defined base categories. It may contain novel class names, attributes, and motions of objects. Incorporating captions in training is proved to be extremely useful in helping improve the models’ scalability.
@@
Moreover, recently, visual language models (VLMs) [16, 17], which pre-train themselves on large-scale image-text pairs, are showing remarkable zero-shot performance on various vision tasks. The VLMs align images and language vocabularies into the same feature space, fulfilling the gap between visual and language data. Many open vocabulary methods effectively eliminate the distinction between close-set and open-set scenarios by utilizing the alignment learned in VLMs, making them highly suitable for practical applications. For example, An open vocabulary object detector can be easily extended to other domains according to the demand without the need to gather relevant data or incur additional labeling expenses.
@@
As open vocabulary models continue to advance rapidly and demonstrate impressive results, it is worthwhile to track and compare recent research on open vocabulary learning. Several surveys work on low-shot learning [18, 19, 20, 21], including few-shot learning and zero-shot learning. There are also several surveys on multi-modal learning [22, 23], including using transformer for vision language tasks [24, 22] and vision language pre-training [25]. However, these works focus on learning with few examples, multi-modal fusion, or pre-training for better feature representation. As far as we know, there haven’t been any surveys that thoroughly summarize the latest developments in open vocabulary learning, including methods, settings, benchmarks, and the use of vision foundation models. Our aim is to fulfill the blank with this work.
@@
Contribution. In this survey, we systematically track and summarize recent literature on open vocabulary learning, including object detection, segmentation, video understanding, and 3D scene understanding. The survey covers the most representative works in each domain by extracting the common technical details. It also contains the background of open vocabulary learning and related concepts comparison, including zero-shot learning (ZSL) [26], open set recognition (OSR) [27, 28], and out-of-distribution detection (OOD) [29]. It also includes large-scale visual language models and representative detection and segmentation works, which makes the survey self-contained. In addition, we present a comprehensive analysis and comparison of benchmarks and settings for each specific domain. As far as we know, we are the first to concentrate on the specific area. Finally, since the field of open vocabulary learning is rapidly evolving, it may not be possible for us to keep up with all the latest developments. We welcome researchers to contact us and share their new findings in this area to keep us updated. Those new works will be included and discussed in the revised version. Our survey project is at https://github.com/jianzongwu/Awesome-Open-Vocabulary.
@@
Survey pipeline. In Sec. 2, we will cover the background knowledge, including definition, datasets, metrics, and related research domains. Then, in Sec. 3, we conduct main reviews on various methods according to different tasks. In particular, we will include the preliminary knowledge of close-set detection and segmentation methods first. Then we present the method details of each direction, including detection, segmentation, video understanding, and 3D scene understanding. In Sec. 4, we compare the results for different tasks and benchmarks. We point out future directions in Sec. 5 and make the conclusions at Sec. 6.
@@
Figure 1: Concepts comparison between open-set/open world/out-of-distribution detection (OOD), zero-shot and open vocabulary. Different shapes represent different novel categories. Colors represent the predictions of the novel objects. (a), in the open-set/Open World/OOD settings, the model only needs to identify novel classes and mark them as ’unknown’. (b) in the zero-shot setting, a model must classify novel classes into specific categories. (c) in the open vocabulary settings, the model can classify novel classes with the help of large language vocabulary knowledge $C_{L}$.
@@ 2 Background
Overview. In this section, we first present the concept definition of open vocabulary and related concepts comparison. Then, we present a historical review of open vocabulary learning and point out several representatives. Next, we present the common datasets and metrics. We also present the unified notations for open vocabulary object detection and segmentation tasks. Finally, we review the related research domains.
Figure 2: Timeline of open vocabulary learning. The gray boxes indicate representative works. Green boxes indicate the foundation models and VLMs. In the field of open vocabulary learning, many works focus on exploiting the knowledge learned by pre-trained vision foundation models like Swin [30] and VLMs like CLIP [16]. Recently, some works also explore the use of diffusion models in this setting.
@@ 2.1 Concept Definition
Taking the classification task as an example, the traditional supervised methods assume that the training data and testing data share the same closed-set label space. However, the model trained under this assumption cannot be extended to new or different categories. To address this issue, researchers have introduced different concepts like open-set learning and zero-shot learning, ultimately leading to open vocabulary learning.
@@
We first compare these closely related concepts in Fig. 1, including open-set, open world, OOD, zero-shot learning, and open vocabulary learning. These concepts have similar settings and notations. In particular, the examples are classified into base and novel classes (or out-of-distribution examples). The base classes are used for training, while the novel classes are not. We denote the base classes in the label set space $C_{B}$ and the novel classes in the label set space $C_{N}$.
The open-set, open world, and OOD tasks share the same testing procedure. Models only need to identify novel categories and set them as one label, named ’unknown’. Models do not need to identify the specific class categories in such settings. The process is shown in Fig. 1(a). Then, on the contrary, zero-shot learning goes beyond these three settings by further carrying out the classification of the novel objects, as shown in Fig. 1 (b). From such a perspective, zero-shot learning is a more challenging setting.
@@
Due to the lower performance on novel classes in practice and the easier acquisition of image-text pairs, recent works [12, 13] propose the open vocabulary settings. It allows the use of additional low-cost training data or pre-trained vision language models like CLIP, which have much larger language vocabularies $C_{L}$. $C_{L}$ may contain the concepts of both $C_{B}$ and $C_{T}$, which enables model to generalize across novel classes. We summarize the concept’s differences as follows:
$\bullet$ Open-Set/Open World/OOD. The three tasks only demand the model to identify unknown objects and reject them from the classification. The model only needs to find unknown classes or examples without the need to classify them into different classes.
$\bullet$ Zero-Shot Learning. The model has strictly trained on $C_{B}$ class examples where the $C_{N}$ is unavailable during training. The model needs to predict the specific categories for novel objects.
$\bullet$ Open Vocabulary Learning. Based on the zero-shot learning setting, one can acquire an additional low-cost training source or knowledge $C_{L}$, which is large-scale language vocabulary. For example, image captions are the most commonly used language data to help extend the semantic space of an open vocabulary model. One can also use the pre-trained vision language models by adapting them for novel class classification. Note that $C_{L}$ is not strictly required to contain $C_{B}$ or $C_{N}$, as the language vocabulary may not cover all the class names in the vision data. On the contrary, $C_{L}$ may also have words out of the pre-defined novel categories, which can further extend models’ generalizability.
@@ 2.2 History and Roadmap
Before introducing the open vocabulary setting, it is necessary to review the progress of open vocabulary learning. In Fig. 2, we summarize the timeline of open vocabulary learning. Localization and classification of arbitrary objects in the wild have been challenging problems due to the limitations of existing datasets. The concept of open vocabulary in scene understanding comes from the work [31], where the authors build a joint image pixel and word concept embedding framework. The concept is hierarchically divided. Then, with the rise of BERT [9] in NLP, multi-modal pre-training was well studied. Motivated by the process of vision language pre-training, OVR-CNN [12] proposed the concept of open vocabulary object detection, where the caption data are used for connecting novel classes semantics and visual region. Later on, CLIP was proposed and open-sourced. After that, VilD [13] is the first work that uses the knowledge of CLIP to build open vocabulary object detection. Meanwhile, LSeg [15] first explored the CLIP knowledge of language-driven segmentation tasks. After these works, recently, there have been more and more works on improving the performance of open vocabulary detectors or building new benchmarks for various settings. Recently, PADing [32] explored the unified open vocabulary learning framework for three different segmentation tasks. SAM [33] is proposed to build the segmentation foundation model, which is trained by billion-level masks. Combing with CLIP, SAM also can achieve good zero-shot segmentation without any fine-tuning. Recently, with the rapid process of large language model (LLM) [34], large open vocabulary learning becomes a more promising direction.

### v4

@@ Abstract
In the field of visual scene understanding, deep neural networks have made impressive advancements in various core tasks like segmentation, tracking, and detection. However, most approaches operate on the close-set assumption, meaning that the model can only identify pre-defined categories that are present in the training set. Recently, open vocabulary settings were proposed due to the rapid progress of vision language pre-training. These new approaches seek to locate and recognize categories beyond the annotated label space. The open vocabulary approach is more general, practical, and effective than weakly supervised and zero-shot settings. This paper thoroughly reviews open vocabulary learning, summarizing and analyzing recent developments in the field. In particular, we begin by juxtaposing open vocabulary learning with analogous concepts such as zero-shot learning, open-set recognition, and out-of-distribution detection. Subsequently, we examine several pertinent tasks within the realms of segmentation and detection, encompassing long-tail problems, few-shot, and zero-shot settings. As a foundation for our method survey, we first elucidate the fundamental principles of detection and segmentation in close-set scenarios. Next, we examine various contexts where open vocabulary learning is employed, pinpointing recurring design elements and central themes. This is followed by a comparative analysis of recent detection and segmentation methodologies in commonly used datasets and benchmarks. Our review culminates with a synthesis of insights, challenges, and discourse on prospective research trajectories. To our knowledge, this constitutes the inaugural exhaustive literature review on open vocabulary learning. We keep tracing related works at https://github.com/jianzongwu/Awesome-Open-Vocabulary.
@@ 1 Introduction
Deep neural networks have revolutionized scene understanding tasks, including object detection, segmentation, and tracking [1, 2, 3, 4]. Nonetheless, using conventional approaches for real-world applications can be challenging due to limitations such as inadequate class annotations, close-set class definitions, and costly labeling expenses. These limitations can increase the difficulty and cost of implementing a deep model on new scenes, particularly when the number of categories or concepts in the scene is much larger than what is included in the training dataset.
@@
For example, object detection is a core computer vision task involving scene understanding. It requires human annotations for each category and each object location, which can be costly and time-consuming. For instance, the COCO dataset [5], widely used for benchmarking object detection algorithms, only includes 80 categories. But in reality, natural scene images often have more than 80 different types of objects. We would need to incur significant annotation costs to extend object detectors to cover all these categories. Current research focuses on developing methods to train more flexible object detectors on the subset of COCO with base classes and let them identify new or unfamiliar objects without requiring additional annotations.
@@
Several previous solutions adopt zero-shot learning (ZSL) [6, 7, 8]. These approaches extend a detector to generalize from annotated (seen) object classes to other (unseen) categories. The annotations of seen object classes are used during the training, while the annotations of unseen classes are strictly unavailable during training. Most approaches adopt word embedding projection to constitute the classifier for unseen class classification.
@@
However, these approaches come with several limitations. ZSL, in particular, is highly restrictive. Typically, these methods lack examples of unseen objects and treat these objects as background objects during training. As a result, during inference, the model identifies novel classes solely based on their pre-defined word embeddings [9, 10], thereby limiting exploration of the visual information and relationships of those unseen classes. This is why ZSL approaches have been shown to yield unsatisfying results in novel classes.
@@
Open vocabulary learning is proposed to handle the above issue and make the detector extendable. It has been successfully applied to multiple tasks, e.g., segmentation, detection, video understanding, scene understanding, etc. In particular, open vocabulary object detection [11, 12] is first proposed. Also, open vocabulary segmentation is proposed [13, 14]. Similarly, the model trains on base classes and inferences on both base and novel classes. The critical difference between zero-shot learning and open vocabulary is that one can use visual-related language vocabulary data like image captions as auxiliary supervision in open vocabulary settings. The motivations to use language data as auxiliary weak supervision are: 1) Language data requires less labeling effort and thus is more cost-friendly. The visual-related language data, like image captions, is widely available and more affordable than box or mask annotations. 2) Language data provides a more extensive vocabulary size and thus is more extendable and general. For example, words in captions are not limited to the pre-defined base categories. It may contain novel class names, attributes, and motions of objects. Incorporating captions in training has been proven to be extremely useful in helping improve the models’ scalability.
@@
Moreover, recently, visual language models (VLMs) [15, 16], which pre-train themselves on large-scale image-text pairs, show remarkable zero-shot performance on various vision tasks. The VLMs align images and language vocabularies into the same feature space, fulfilling the visual and language data gap. Many open vocabulary methods effectively eliminate the distinction between close-set and open-set scenarios by utilizing the alignment learned in VLMs, making them highly suitable for practical applications. For example, An open vocabulary object detector can be easily extended to other domains according to the demand without the need to gather relevant data or incur additional labeling expenses.
@@
As open vocabulary models continue to advance rapidly and demonstrate impressive results, it is worthwhile to track and compare recent research on open vocabulary learning. Several surveys work on low-shot learning [17, 18, 19, 20, 21], including few-shot learning and zero-shot learning. There are also several surveys on multi-modal learning [22, 23], including using transformer for vision language tasks [24, 22] and vision language pre-training [25]. However, these works focus on learning with few examples, multi-modal fusion, or pre-training for better feature representation. As far as we know, there haven’t been any surveys that thoroughly summarize the latest developments in open vocabulary learning, including methods, settings, benchmarks, and the use of vision foundation models. We aim to fulfill the blank with this work.
@@
Contribution. In this survey, we systematically track and summarize recent literature on open vocabulary learning, including object detection, segmentation, video understanding, and 3D scene understanding. The survey covers the most representative works in each domain by extracting the common technical details. It also contains the background of open vocabulary learning and related concepts comparison, including zero-shot learning (ZSL) [26], open set recognition (OSR) [27, 28], and out-of-distribution detection (OOD) [29]. It also includes large-scale visual language models and representative detection and segmentation works, which makes the survey self-contained. In addition, we present a comprehensive analysis and comparison of benchmarks and settings for each specific domain. As far as we know, we are the first to concentrate on the specific area. Finally, since the field of open vocabulary learning is rapidly evolving, we may not be able to keep up with all the latest developments. We welcome researchers to contact us and share their new findings in this area to keep us updated. Those new works will be included and discussed in the revised version.
@@
Survey pipeline. In Sec. 2, we will cover the background knowledge, including definition, datasets, metrics, and related research domains. Then, in Sec. 3, we conduct main reviews on various methods according to different tasks. In particular, we will first include the preliminary knowledge of close-set detection and segmentation methods. Then, we present the method details of each direction, including detection, segmentation, video understanding, and 3D scene understanding. Next, we point out challenges and future directions in Sec. 4 and conclude Sec. 5. Finally, the appendix compares the results for different tasks and benchmarks.
@@
Figure 1: Concepts comparison between open-set/open world/out-of-distribution detection (OOD), zero-shot and open vocabulary. Different shapes represent different novel categories. Colors represent the predictions of the novel objects. (a), in the open-set/open World/OOD settings, the model only needs to identify novel classes and mark them as “unknown”. (b), in the zero-shot setting, a model must classify unknown classes into specific categories. (c), in the open vocabulary settings, the model can classify novel classes with the help of large language vocabulary knowledge $C_{L}$.
@@ 2 Background
Overview. In this section, we first present the concept definition of open vocabulary and related concepts comparison. Then, we present a historical review of open vocabulary learning and point out several representatives. Next, we present the standard datasets and metrics. We also present the unified notations for open vocabulary object detection and segmentation tasks. Finally, we review the related research domains.
Figure 2: Timeline of open vocabulary learning. The gray boxes indicate representative works. Green boxes indicate the foundation models and VLMs. In open vocabulary learning, many works exploit the knowledge learned by pre-trained vision foundation models like Swin [30] and VLMs like CLIP [15]. Recently, some works also explore the use of diffusion models in this setting.
(a) Number of Research Works
(b) Ratio of Different Directions
(c) Different Directions per Year
Figure 3: Summarization on open vocabulary learning works. (a) The number of research works still increases per year. (b) Detection and segmentation have more papers than 3D and video. (c) indicates each direction per year. The results are obtained on 2024/1/15.
@@ 2.1 Concept Definition
We take the classification task for concept illustration. The supervised methods assume that the training data and testing data share the same closed-set label space. However, the model trained under this assumption cannot be extended to new categories. To address this issue, researchers have introduced several new concepts like open-set learning and zero-shot learning, ultimately leading to open vocabulary learning.
@@
These concepts have similar settings and notations. In particular, the examples are classified into base and novel classes (or called out-of-distribution examples). The base classes can be accessed for training, while the novel classes are not. We denote the base classes in the label set space $C_{B}$ and the novel classes in the label set space $C_{N}$.
@@
Due to the low performance on novel classes in earlier works for open-set, open world, and OOD tasks [31, 32, 29, 11] and the easier acquisition of image-text pairs, recent works [11, 12] propose the open vocabulary setting. It allows using additional low-cost training data [33] or pre-trained vision language models like CLIP [15], which have much larger language vocabularies $C_{L}$. It may contain the concepts of both $C_{B}$ and $C_{N}$, but its main goal is to enable the model to generalize across more classes in the open domain. In mathematical terms and data view, open vocabulary learning can be defined as follows:
Training Data ($D_{train}$): The training dataset is a collection of data-label pairs, where each pair consists of an input $x_{i}$ and its associated label $y_{i}$. The input $x_{i}$ can be an image or a video according to the task type, and in scene understanding tasks such as detection and segmentation, the label $y_{i}$ also includes visual labels, such as bounding boxes or masks, in addition to the class labels. Alongside the standard input-label pairs, open vocabulary learning includes vision-aware language vocabulary data, represented as $l_{i}$. The language vocabulary data can be image-caption data or vision-aware class name embeddings in VLMs like CLIP [15]. The language augmented training set can be denoted as $D_{train}=\{(x_{1},y_{1},l_{1}),(x_{2},y_{2},l_{2}),...,(x_{n},y_{n},l_{n})\}$, where $n$ is the training data length, $x_{i}$ is the visual data, $y_{i}$ is the label from base classes $C_{B}$, and $l_{i}$ is the associated language data from a large vocabulary space $C_{L}$. Note that $C_{L}$ is not strictly required to contain $C_{B}$ or $C_{N}$, as the language vocabulary may not cover all the class names in the vision data. On the contrary, $C_{L}$ may also have words out of the pre-defined novel categories, which can further extend models’ generalizability.
Evaluation Data ($D_{eval}$): The evaluation dataset is similarly a collection of data pairs. However, the labels for evaluation data include both base classes and novel classes. This can be represented as $D_{eval}=\{(x^{\prime}_{1},y^{\prime}_{1}),(x^{\prime}_{2},y^{\prime}_{2}),...,(x^{\prime}_{m},y^{\prime}_{m})\}$, where $m$ is the evaluation data length, $y^{\prime}_{i}$ belongs to either $C_{B}$ or $C_{N}$. During the evaluation, open vocabulary methods need to predict $y^{\prime}_{i}$ given $x^{\prime}_{i}$ in the realm $y^{\prime}_{i}\in(C_{B}\cup C_{N})$.
Here, we emphasize the concept differences between open vocabulary learning and analogous concepts as follows:
$\bullet$ Open-Set Learning. Open-set learning aims to classify known classes and reject unknown classes during testing [31, 34, 35]. Concretely, the training data is $D_{train}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{n},y_{n})\}$, where $y_{i}$ is from base classes $C_{B}$. These are also called known classes in open-set learning [31, 36, 37]. The evaluation data is $D_{eval}=\{(x^{\prime}_{1},y^{\prime}_{1}),(x^{\prime}_{2},y^{\prime}_{2}),...,(x^{\prime}_{m},y^{\prime}_{m})\}$, where $y^{\prime}_{i}\in(C_{B}\cup\{u\})$. $u$ is a single class that represents the ‘unknown’ class. Open-set learning tasks do not require further classifying the unknown classes.
$\bullet$ Open World Learning. Open world learning addresses real-world environments’ dynamic and ever-evolving nature by recognizing and learning new categories incrementally over time, without the need for complete system retraining [38, 39]. This process includes classifying known objects, identifying unknowns, labeling the unknowns by humans, and incrementally learning new categories as they are labeled and added to the system. Suppose an open world learning process contains $T$ learning steps. In the $t$-th learning step, where $t\in\{1,2,...,T\}$, the training data is $D_{train}^{t}=\{(x_{1}^{t},y_{1}^{t}),(x_{2}^{t},y_{2}^{t}),...,(x_{n_{t}}^{t},y_{n_{t}}^{t})\}$, where $y_{i}^{t}$ is in the labels annotated in the $t$-th step $C^{t}_{label}$, and $x_{i}^{t}$ is the input data corresponding to the newly added labels. The evaluation data is $D_{eval}^{t}=\{(x^{\prime}_{1},y^{\prime t}_{1}),(x^{\prime}_{2},y^{\prime t}_{2}),...,(x^{\prime}_{m},y^{\prime t}_{m})\}$, where $y^{\prime t}_{i}\in(C^{t}_{B}\cup\{u\})$. $u$ is a single class that represents the ‘unknown’ class. $C^{t}_{B}$ is the known class label set for timestamp $t$. It is an accumulation of labeled classes in the current and previous steps. $C^{t}_{B}=\bigcup_{k=0}^{t}C^{k}_{label}$. Note that evaluation input $x^{\prime}_{i}$ keeps the same across learning steps.
$\bullet$ Out-of-Distribution Detection. Out-of-distribution (OOD) detection focuses on the ability to detect data that is different in some way from the data used during training [29]. The training data is $D_{train}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{n},y_{n})\}$ and is supposed to be sampled from a probability distribution $P_{train}(X,Y)$. During testing, the model encounters data $x^{\prime}_{i}$ that are sampled from a different distribution $P_{OOD}(X^{\prime})$, which is not represented in the training dataset. The goal is to identify or appropriately handle these OOD samples. Metrics for OOD detection often involve measuring the model’s certainty or confidence in its predictions, using scores like softmax probability, and checking whether it correctly identifies OOD samples [29].
$\bullet$ Zero-Shot Learning. Zero-shot learning aims to recognize objects or concepts not seen during training. The training data is $D_{train}=\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{n},y_{n})\}$, where y is from base classes $C_{B}$, which is usually called seen classes in zero-shot learning [40, 26]. The evaluation data is $D_{eval}=\{(x^{\prime}_{1},y^{\prime}_{1}),(x^{\prime}_{2},y^{\prime}_{2}),...,(x^{\prime}_{m},y^{\prime}_{m})\}$, where $y^{\prime}_{i}\in C_{N}$. $C_{N}$ is novel classes or unseen classes. Zero-shot learning tasks require models to make clear classifications among the new unseen classes.
We briefly compare these concepts in Fig. 1, including open vocabulary, open-set, open world, OOD, and zero-shot.
@@ 2.2 History and Roadmap
Before introducing the open vocabulary setting, reviewing the progress of open vocabulary learning is necessary. In Fig. 2, we summarize the timeline of open vocabulary learning. Localization and classification of arbitrary objects in the wild have been challenging problems due to the limitations of existing datasets. The concept of open vocabulary in scene understanding comes from the work [41], where the authors build a joint image pixel and word concept embedding framework. The concept is hierarchically divided. Then, multi-modal pre-training was well studied with the rise of BERT [9] in NLP. Motivated by the process of vision language pre-training, OVR-CNN [11] proposed the concept of open vocabulary object detection, where the caption data are used for connecting novel classes semantics and visual region. Later on, CLIP was presented and open-sourced. After that, VilD [12] is the first work that uses the knowledge of CLIP to build open vocabulary object detection. Meanwhile, LSeg [14] first explored the CLIP knowledge of language-driven segmentation tasks. After these works, recently, there have been more and more works on improving the performance of open vocabulary detectors or building new benchmarks for various settings. SAM [42] is proposed to build the segmentation foundation model, which is trained by billion-level masks. Combined with CLIP, SAM can also achieve good zero-shot segmentation without fine-tuning. Recently, with the rapid process of large language model (LLM) [43], open vocabulary learning has become a more promising direction since more language knowledge can be embedded in multi-modal architecture. As shown in Fig. 3(a), the number of research works in open vocabulary learning has increased significantly since 2021. We also summarize the statistics of different directions in Fig. 3(b) and (c). The details of these directions can be found in Sec. 3.